A certain weak base has a Kt of 8.60 x 10-7. What concentration of this base will produce a pH of 10.317 Number
Answer #1

Let the concentration of B c molar

use:

pH = -log [H+]

10.31 = -log [H+]

[H+] = 4.898*10^-11 M

use:

[OH-] = Kw/[H+]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[OH-] = (1.0*10^-14)/[H+]

[OH-] = (1.0*10^-14)/(4.898*10^-11)

[OH-] = 2.042*10^-4 M

B dissociates as:

B +H2O -----> BH+ + OH-

c 0 0

c-x x x

Kb = [BH+][OH-]/[B]

Kb = x*x/(c-x)

8.6*10^-7 = 2.042*10^-4*2.042*10^-4/(c-2.042*10^-4)

c-2.042*10^-4 = 4.847*10^-2

c=4.868*10^-2 M

Answer: 4.87*10^-2 M